OpenStudy (anonymous):

Can sombody please help me simplify this expression.? -5+i over 2i

OpenStudy (anonymous):

\(\large \frac{-5+i}{2i} \) \(\large \frac{-5+i}{2i}\cdot \color {red}{ \frac{2i}{2i}} \) can you multiply the top first?

OpenStudy (anonymous):

ummmm, is it -10+2i over 4i??

OpenStudy (anonymous):

I'm sorry, I don't really know how to do this/:

OpenStudy (anonymous):

\(\large \frac{-5+i}{2i}\cdot \color {red}{ \frac{2i}{2i}} \) = \(\large \frac{\color {red}{2i}(-5+i)}{\color {red}{2i}\cdot 2i} \) = \(\large \frac{\color {red}{2i}(-5)+\color {red}{2i}(i)}{4i^2} \) can you finish?

OpenStudy (anonymous):

-10i+2i over 4i^2 ?

OpenStudy (anonymous):

@dpaInc where'd you go?

OpenStudy (anonymous):

you should have -10i + 2i^2 over 4i^2.... now make a substitution using the fact that i^2 = -1...

OpenStudy (anonymous):

so, the answer is -12i over -4......?

OpenStudy (anonymous):

@dpaInc ?

OpenStudy (anonymous):

by "simplify" you mean write in standard form as \(a+bi\)?

OpenStudy (anonymous):

Idk/: The question just says "Simplify the expression"

OpenStudy (anonymous):

multiply top and bottom by \(i\) and recall that \(i^2=-1\) that gives you \[\frac{-5+i}{2i}\times \frac{i}{i}=\frac{-5i+i^2}{2i^2}=\frac{-5i-1}{-2}\]

OpenStudy (anonymous):

now you can rewrite as \[\frac{1}{2}-\frac{5}{2}i\] if you like

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" it is the lazy math teacher way of saying "give me the answer i want" there is however standard form of a complex number, and that is \(a+bi\) where \(a\) and \(b\) are real numbers

OpenStudy (anonymous):

actually i made a mistake above , answer should be \[\frac{1}{2}+\frac{5}{2}i\]

OpenStudy (anonymous):

Ok thank you so much! That helped me understand it a whole lot better.

OpenStudy (anonymous):

yw